Here is the ledger for Sandhill Co..

Cash

Oct. 1

8,810

Oct. 4

2,210

Oct. 10

2,790

Oct. 12

3,310

Oct. 10

9,810

Oct. 15

400

Oct. 20

850

Oct. 30

450

Oct. 25

3,810

Oct. 31

650

Accounts Receivable

Oct. 6

950

Oct. 20

850

Oct. 20

1,070

Supplies

Oct. 4

2,210

Oct. 31

1,990

Equipment

Oct. 3

4,810

Notes Payable

Oct. 10

9,810

Accounts Payable

Oct. 12

3,310

Oct. 3

4,810

Common Stock

Oct. 1

8,810

Oct. 25

3,810

Dividends

Oct. 30

450

Service Revenue

Oct. 6

950

Oct. 10

2,790

Oct. 20

1,070

Salaries and Wages Expense

Oct. 31

650

Supplies Expense

Oct. 31

1,990

Rent Expense

Oct. 15

400


(a)

Reproduce the journal entries for only the transactions that occurred on October 1, 10, and 20. (If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually. Record journal entries in the order presented in the problem.)

Date

Account Titles and Explanation

Debit

Credit

choose a transaction date                                                          Oct. 1Oct. 10Oct. 20

enter an account title for issued stock for cash

enter a debit amount

enter a credit amount

enter an account title for issued stock for cash

enter a debit amount

enter a credit amount

(Issued stock for cash)

choose a transaction date                                                          Oct. 1Oct. 10Oct. 20

enter an account title for received cash for services provided

enter a debit amount

enter a credit amount

enter an account title for received cash for services provided

enter a debit amount

enter a credit amount

(Received cash for services provided)

choose a transaction date                                                          Oct. 1Oct. 10Oct. 20

enter an account title for issued note payable for cash

enter a debit amount

enter a credit amount

enter an account title for issued note payable for cash

enter a debit amount

enter a credit amount

(Issued note payable for cash)

choose a transaction date                                                          Oct. 1Oct. 10Oct. 20

enter an account title for received cash in payment of account

enter a debit amount

enter a credit amount

enter an account title for received cash in payment of account

enter a debit amount

enter a credit amount

(Received cash in payment of account)

choose a transaction date                                                          Oct. 1Oct. 10Oct. 20

enter an account title for billed clients for services provided

enter a debit amount

enter a credit amount

enter an account title for billed clients for services provided

enter a debit amount

enter a credit amount

(Billed clients for services provided)


(b)

Prepare a trial balance at October 31, 2022.

SANDHILL CO.
Trial Balance

choose an accounting period                                                          For the Month Ended October 31, 2022For the Year Ended October 31, 2022October 31, 2022

Debit

Credit

enter an account title

$enter a debit balance

$enter a credit balance

enter an account title

enter a debit balance

enter a credit balance

enter an account title

enter a debit balance

enter a credit balance

enter an account title

enter a debit balance

enter a credit balance

enter an account title

enter a debit balance

enter a credit balance

enter an account title

enter a debit balance

enter a credit balance

enter an account title

enter a debit balance

enter a credit balance

enter an account title

enter a debit balance

enter a credit balance

enter an account title

enter a debit balance

enter a credit balance

enter an account title

enter a debit balance

enter a credit balance

enter an account title

enter a debit balance

enter a credit balance

enter an account title

enter a debit balance

enter a credit balance

$enter a total for the debit column

$enter a total for the credit column
Answer #1

On looking at the ledgers, observe and match for corresponding amounts:- for example, on oct 1, observe cash ledger, it is de

1,070 20-Oct Accounts Receivable Service Revenue (to record service revenue received on account) 1,070

SANDHILL CO. Trial Balance for the month ended October 31,2022 Debit Credit Cash 19,050 Accounts receivable 1,170 Supplies 22

Trial balance figures are the net of credit and debit figures of each account. For Example:

Cash Balance =8810+2790+9810+850+3810-2210-3310-400-450-650 = 19,050. Debit

Accounts receivable = 950+1070-850 = 1,170 Debit

Supplies = 2210-1990 = 220 debit

Accounts payable =4810-3310 = 1,500 credit

Common Stock =8810+3810 = 12,620 credit

Service Revenues =950+2790+1070 = 4,810 credit

Hope this solves your query. Feel free to ask any further query in the comment section.